# WildGraph: Realistic Graph-based Trajectory Generation for Wildlife

### v1

@@ 3.3. H3 Grid System
As mentioned above, we utilize the H3 geospatial indexing system. H3 is an open-source library for creating and managing hierarchical spatial indexes. The core concept of H3 is the use of hexagonal grids that partition the map into smaller spatial units, at different zoom levels. The developers cite various advantages that include: (1) capacity to model large spatial regions and have better descriptive ability of spherical surfaces; (2) provides a single string-based address (e.g. 821fa7fffffffff) that encompasses a geographic region of points; (3) built-in hierarchical description of a region at 14 zoom levels (Uber Technologies, Inc., cess).
@@ 4. Methodology
Figure 3. WildGraph Framework
@@ 4.1. Hierarchical Network Generator
(a) *
[.33]Refer to caption [.66]Refer to caption
@@
Note that excessive zooming results in a large network that consists of too many nodes, which may limit the generalization of the model, and increase its complexity. The optimal threshold $r$ is learned from the training data. It depends on properties of the dataset, and may be influenced by the resulting average density of the graph. In the experimental section, we empirically investigate how the choice of $r$ affects the evaluation metrics.
@@
1:  Initialize $zoom\leftarrow 2$ //Starting H3 regions at this zoom level
2:  Initialize $P\leftarrow(p_{1},\ldots,p_{i\times m})$
3:  $H\leftarrow Assign\_H3(P)$ //Assign H3 address to each point
4:  while $\forall h3\in H3\mid diameter(h3)>r$ do
5:     $zoom\leftarrow zoom+1$
6:     $P_{s}\leftarrow P[h3]$ //Filter points that need granular H3 regions
7:     $H_{s}\leftarrow Assign\_H3(P_{s})$
8:     Update $P$ with $P_{s}$ //Propagate updates
9:     Update $H$ with $H_{s}$ //Propagate updates
10:  end while
11:  return  $P,H$
Algorithm 1 Regionalize Trajectories([$T_{1},\dots,T_{n}$], m)
@@ 4.2. Network Graph Embedding
Given, the long-horizon nature of our trajectories, we select a higher value of $q$ to explore more distant nodes. The resulting embeddings learned by node2Vec encode rich structural information about nodes. Node embeddings also significantly reduce the dimensionality of the problem and facilitate out of policy exploration: transitions that are not observed in the real dataset, but are potentially realistic transitions.
@@ 4.3. Variational Recurrent Network
The Variational Recurrent Network (VRN) uses the embedded sequences of trajectories to train the model by processing them sequentially. It combines elements of recurrent neural networks (RNNs) with variational inference techniques to model the sequential decision-making under uncertainty. Analogous to conventional RNNs, the VRN incorporates recurrent connections that allow it to maintain a hidden state representing the network’s memory of past observations and actions. This hidden state evolves over time as the network processes the input sequentially.
The VRN is trained using a combination of unsupervised learning and recurrence techniques. Unsupervised learning is used to train the network to predict the next observation or action given the current observation and hidden state. The utilized observation is the current belief vector over the states, obtained as the raw output of the VRN. Given the recurrence nature of VRN, the belief vector is trained on the history of the previous states. This approach may be viewed as learning a transition kernel from the current state to future states while considering the history of the trajectory. Post-training, the hidden state, the output of the bottleneck layer in the autoencoder, is statistically treated to reflect the possible transitions per the training trajectories, though it is represented in a compressed space.
The VRN utilizes an encoder-decoder model. Given a pair of region-mapped and graph embedded points $(h3_{i},h3_{i+1})$, the encoder maps $h3_{i+1}$ to the latent space representation $z_{i+1}$. The decoder attempts to reconstruct the input vector $h3_{i+1}$ from a concatenation of $z_{i+1}$ with $h3_{i}$, modified with positional encoding. The converged $z_{i+1}$ after training is captured and added to the multi-dictionary of $h3_{i}$.
During inference, the VRN makes sequential decisions by iteratively updating its hidden state based on observed inputs and selecting actions according to its learned policy. The network’s belief state is updated, allowing it to maintain a probabilistic estimate of the environment’s true state given the past states.
@@ 4.3.1. Positional Encoding
Positional encoding (PE) assigns unique representations to each position in the sequence. This allows the model to differentiate between regions based on their position in the sequence. For example, a bird may remain in a given region for several days before proceeding to the next region. The PE enables the model to learn from the sequential context effectively. We use a simple approach by concatenating the raw index of the position to the input of the decoder. This helps the network to associate certain transitions with the position, while avoiding the risk that the VRN may fail to learn a more sophisticated method given the small data nature of the problem.
@@ 4.4. Latent Dictionary
After training, the representation of each possible state $h3$ is passed through the encoder to find the corresponding latent space realization $z$. Using the frequentist approach, for each node $h3$, the possible transitions are collected in a dictionary and then replaced by their corresponding latent space realizations. When a certain node appears as the current state in our VRN, we sample uniformly from the latent dictionary of this node. This is akin to sampling the stationary next state based on the possible transitions in the training set, but in latent space. However, these stationary transitions are combined with the belief vector, thus learning the non-stationary transitions, while taking into consideration all previously generated states.
@@
For all pairs of transitions between nodes in the training set: $(k_{i},k_{j})$, the dictionary is updated: $D^{\prime}=D\cup\{k_{i}:(l_{j})\}$ where $l_{j}=enc(k_{j})$, and $enc$ the trained encoder from the VRN.
@@ 4.5. Occupancy Sampler
The occupancy sampler chooses a dot (a very small region) and then from the dot, it chooses a point at random. The sampling is conducted based on the heatmap of the region as depicted in Figure 5. The sampler proceeds as follows:
@@
    Split the region into dots, which in our case are H3 regions with a high zoom that fall within the region.
@@
The intuition behind using a two-step hierarchical sampler (i.e., going to dots and then choosing points) as opposed to choosing points from a heatmap directly is as follows. If we represented a region as a set of points and used a heatmap to choose the points directly, then we would be choosing points from existing trajectories with high probability. By introducing a small region that we select based on the heatmap and randomizing the choice of the point from the dot, we introduce a level of generalization that is useful in introducing additional stochasticity in point selection. An added advantage of this approach is it helps ensure the points selected have geographic validity, i.e., are close enough to actual points because of how we choose the output points based on a dot, which is a very small region. This can been visually confirmed from the generated paths in Figure 6(a), where WildGraph trajectories, much like real trajectories, almost always avoiding congregating in the middle of the sea, while this is not the case with some of the benchmark methods, even competitive methods such as WildGEN.
Figure 5. Occupancy Sampling of dots based on heatmap
@@ 5.1. Baselines
  • Generative Adversarial Network (GAN)
  • Transformers
@@ 5.1.2. WildGEN
  • Minimum Bounding Region: drops any trajectories that venture out of a minimum bounding polygon of the set of known points.
5.1.3. GANs
Generative Adversarial Network (GAN) is a model that consist of two neural networks, the generator and the discriminator, which engage in a minimax game. The generator aims to produce realistic synthetic data samples from random noise, while the discriminator learns to distinguish between real and generated samples. Through adversarial training, GANs progressively improve both the generator’s ability to create realistic data and the discriminator’s proficiency in discerning real from fake. This adversarial process drives the generator to generate increasingly realistic samples (Goodfellow et al., 2020).
@@ 5.1.4. Transformer
Transformers are networks that have the capacity to auto-regressively generate sequences of tokens based on attending to previous tokens. The HNG facilitates the use of H3 word-based language transformers, such that a trajectory is a sentence that consists of $m$ words: each of which is an H3 address.
@@ 5.2. Evaluation Metrics
In order to measure Path Similarity, we use three distance metrics: Hausdorff Distance, which measures the steepest gap, and Dynamic Time Warping (DTW) which measures the spatial similarity over the whole path. In addition, we also measure the Final Displacement Error (FDE) which we define as the gap between the final positions of a generated trajectory and the nearest real trajectory: $\text{FDE}=\|p_{t}^{m}-\hat{p}_{s}^{m}\|$.
On the other hand, we measure Likeness using Pearson Correlation Coefficient (r-Coeff) and Chi-squared ($chi^{2}$) similarity, between the generated trajectories and the held-out real trajectory set.
For each generated trajectory, we compare it to every trajectory from the test set, and use the closest match. The reported distance metric for the experiment is the mean of these values over a fixed number of generated trajectories.
5.2.1. Coverage
In addition, we introduce a new evaluation metric called coverage (cov) to measure how representative the generated trajectories are of the test set of real trajectories. This measure, which is captured for each distance evaluation metric above, quantifies the percentage of the held-out real data that provided the measurement from the generated trajectories. For example, given generated trajectory $S_{i}$, if the best Hausdorff or DTW measure was observed when compared to a real trajectory $T_{j}$, we add $T_{j}$ in the coverage list. This measure quantifies the fraction of trajectories added in the coverage list of each distance metric.
@@ 5.3. Datasets
We utilize two dataset sets of migratory birds obtained from Movebank [^2] https://www.movebank.org. Both datasets were processed using Moveapps (Kölzsch et al., 2022) to limit observations to one per day for each subject. For both datasets, any subject that had significant missing readings were dropped. Subjects that were tracked over multiple years were split into multiple samples. Simple data interpolation was utilized to fill in any remaining isolated gaps in the trajectories. The two datasets are:
  • The Western Palearctic greater white-fronted geese dataset (geese dataset) (Kölzsch et al., 2019), which was clipped to a window between March 1 to September 1 (185 days) to capture northward migration. The resultant dataset consists of 60 trajectories over a period of 185 days.
  • The White Stork Ciconia ciconia dataset (stork dataset) (Carlson et al., 2021), which was clipped to a window between August 15 to November 1 (79 days) to capture southward migration. The resultant dataset consists of 77 trajectories over a period of 79 days.
@@ 5.4. Experimental Settings
Preliminary tests were performed to evaluate the convergence properties of each method on a fixed learning rate of $31e-$ 3 using Adam optimizer. For the GAN experiment, additional decay rate parameters were used to induce convergence.
Each experiment used 5-fold cross validation to generalize given the small data nature of the problem. All reported values are the average of 5 identical experiments. This choice was made to normalize the randomness of the generation process, and report more consistent results.
The experiments were conducted utilizing a single 80GB A100 Nvidia GPU paired with 8 CPUs, operating on Python 3.8, PyTorch 2.1.1, and CUDA 11.8.
@@ 5.4.1. Distributional Settings
In order to measure the likeness (r-Coeff and $chi^{2}$), we utilize the discrete cluster distributions between the generated points and the test set. In order to create the clusters, the optimal cluster size for K-means was pre-calculated and used as input in the distributional likeness measurements. The cluster sizes were determined using a Silhouette graph and selected visually. Specifically, the cluster sizes for the geese and stork datasets are 16 and 14, respectively.

### v2

@@ 3.3. H3 Grid System
As mentioned above, we utilize the H3 geospatial indexing system. H3 is an open-source library for creating and managing hierarchical spatial indexes. The core concept of H3 is the use of hexagonal grids that partition the map into smaller spatial units at different zoom levels. The developers of H3 cite various advantages that include: (1) capacity to model large spatial regions and have better descriptive ability of spherical surfaces; (2) provides a single string-based address (e.g. 821fa7fffffffff) that encompasses a geographic region of points; (3) built-in hierarchical description of a region at 14 zoom levels (Uber Technologies, Inc., cess).
@@ 4. Methodology
(a) Training
(b) Generation
Figure 3. WildGraph Framework, including both the training and generation steps. Some details are not depicted for clarity.
@@ 4.1. Hierarchical Network Generator
[.33]Refer to caption [.6]Refer to caption
@@
Note that excessive zooming results in a large network that consists of too many nodes, which may limit the generalization of the model, and increase its complexity. The optimal threshold $r$ depends on properties of the dataset, and may be influenced by the resulting average density of the graph. In the experimental section, we empirically investigate how the choice of $r$ affects the evaluation metrics.
@@
1:  $zoom\leftarrow 2$Initialize H3 regions at this zoom level}
2:  $P\leftarrow Get\_Points(T_{train})$Get set of train-set trajectories}
13:  $H\leftarrow Assign\_H3(P,zoom)$Assign H3 to each point}
4:  $H3\leftarrow SET(H)$Get unique H3 regions}
5:  for $h3\in H3$ do
6:     if $diameter(h3)>r$ then
7:        $zoom\leftarrow zoom+1$
8:        $P_{s}\leftarrow P[h3]$Filter points from $P$ in $h3$}
9:        $H_{s}\leftarrow Assign\_H3(P_{s},zoom)$
10:        $H3\leftarrow H3\setminus\{h3\}$Remove $h3$ from $H3$}
11:        $H3\leftarrow H3\cup H_{s}$Add new $H_{s}$ to $H3$}
12:     end if
213:  end for
14:  return  $H3$
Algorithm 1 Regionalize Trajectories($T_{train}$, r)
@@ 4.2. Network Graph Embedding
Given, the long-horizon nature of our trajectories, we select a higher value of $q$ to explore more distant nodes. The resulting embeddings learned by node2Vec encode rich structural information about nodes. Node embeddings also significantly reduce the dimensionality of the problem and facilitate off-policy exploration: transitions that are not observed in the real dataset, but are potentially realistic transitions.
@@ 4.3. Variational Recurrent Network
The VRN uses the embedded sequences of trajectories to train the model by processing them sequentially. It combines elements of recurrent neural networks (RNNs) with variational inference techniques to model the sequential decision-making under uncertainty. Analogous to conventional RNNs, the VRN incorporates recurrent connections that allow it to maintain a hidden state representing the network’s memory of past observations and actions. This hidden state evolves over time as the network processes the input sequentially.
The VRN is trained using a combination of self-supervised learning and recurrence techniques. Self-supervised learning is used to train the network to predict the next observation or action given the current observation and hidden state using an encoder-decoder model. Given a pair of region-mapped and graph embedded points $(h3_{emb_{i}},h3_{emb_{i+1}})$, the encoder maps $h3_{emb_{i+1}}$ to the latent space representation $z_{i+1}$. The decoder attempts to reconstruct the input vector $h3_{emb_{i+1}}$ from a concatenation of $z_{i+1}$ with $h3_{emb_{i}}$, modified with positional encoding. The converged $z_{i+1}$ after training is captured and added to the multi-dictionary of $h3_{i}$.
4.3.1. Model Formulation
For each transition $(h3_{i},h3_{i+1})$ observed in any given trajectory in $\mathcal{G}$, The input embedding $h3_{emb_{i}}$ is computed as:
$h3_{emb_{i}}=\text{Embedding}(h3_{i})$
The latent representation $z_{i}$ is obtained using a MLP encoder ($\theta$’s are learnt weights):
$z_{i+1}=\text{MLP}_{\text{enc}}(h3_{emb_{i+1}};\theta_{\text{enc}})$
The decoder’s hidden state $h_{\text{dec}}$ is computed using a MLP decoder:
$h_{\text{dec}}=\text{MLP}_{\text{dec}}(h3_{emb_{i}},z_{i+1},PE;\theta_{\text{dec}})$
The logits are computed using a linear transformation of the decoder’s output:
$\mathbf{o}=\text{Linear}(\mathbf{h}_{\text{dec}},\mathbf{w}_{\text{out}},\mathbf{b}_{\text{out}})$
where $\mathbf{w}_{\text{out}}$ is the weight matrix and $\mathbf{b}_{\text{out}}$ is the bias vector.
The logits are transformed into probabilities using the softmax function:
$prob_{k}=\frac{\exp(o_{k})}{\sum_{j=1}^{V_{\text{out}}}\exp(o_{j})}$
where $prob_{k}$ is the predicted probability for the $k$-th class and $V_{\text{out}}$ is the size of the output vocabulary, i.e. the number of regions.
4.3.2. Loss Function
The cross-entropy loss between the true label $h3_{i+1}$ and the predicted probabilities $\mathbf{prob}$ is computed as:
$\mathcal{L}=-\sum_{k=1}^{V_{\text{out}}}enc(h3_{i+1})\log(prob_{k})$
where $enc(h3_{i+1})$ is the one-hot encoded true label.
Post-training, the hidden state (the output of the bottleneck layer in the autoencoder) reflects the possible transitions learned from the training trajectories, though it is represented in a compressed space.
4.3.3. Generation
During generation, we first sample the latent space, add positional encodings (PE), and append the belief vector, which is a weighted sum of the top-$y$ states based on their probabilities ($\sum_{i=1}^{y}\text{Prob}(h3_{i})\times h3_{emb_{i}}$), at the previous time step. This belief vector considers all previously visited states autoregressively, rendering each generated trajectory unique. Empirical results demonstrate that even a set of 1000 generated trajectories exhibits a unique ordering of regions. We formally prove the low likelihood of duplicate generated trajectories in Appendix A.
@@ 4.3.4. Positional Encoding
Positional encoding (PE) assigns unique representations to each position in the sequence. This allows the model to differentiate between regions based on their position in the sequence. For example, a bird may remain in a given region for several days before proceeding to the next region. PE enables the model to learn from the sequential context effectively. We use a simple approach by concatenating the raw index of the position to the input of the decoder. This helps the network to associate certain transitions with the position, while avoiding the risk that the VRN may fail to learn a more sophisticated method given the small data nature of the problem.
@@ 4.4. Latent Dictionary
After training, the representation of each possible state $h3$ is passed through the encoder to obtain the corresponding latent space realization $z$. Using the frequentist approach, for each node $h3$, the possible transitions are collected in a dictionary and then replaced with their corresponding latent space realizations. When a certain node appears as the current state in our VRN, we sample uniformly from the latent dictionary of this node. This is akin to sampling the stationary next state based on the possible transitions in the training set, but in latent space. However, these stationary transitions are combined with the belief vector, thus learning the non-stationary transitions, while taking into consideration all previously generated states. As discussed, this belief vector is computed based on the $y$ most probable $h3$ regions and their corresponding node2vec embeddings.
@@
For all pairs of transitions between nodes in the training set: $(k_{i},k_{j})$, the dictionary is updated: $D^{\prime}=D\cup\{k_{i}:(l_{j})\}$ where $l_{j}=MLP_{enc}(k_{j})$.
@@ 4.5. Occupancy Sampler
The occupancy sampler chooses a dot (a very small region) and then from the dot, it chooses a point at random. The sampling is conducted based on the heatmap of the region as depicted in Figure 5. The process is as follows:
@@
    Divide the region into dots, where each dot represents an H3 region with a high zoom level that lies within the region.
@@
The intuition behind using a two-step hierarchical sampler (i.e., going to dots and then choosing points) as opposed to choosing points directly from the heatmap is as follows. If we represented a region as a set of points and used a heatmap to choose the points directly, then we would be choosing points from existing trajectories with high probability. By introducing a small region that we select based on the heatmap and randomizing the choice of the point from the dot, we introduce a level of generalization that is useful in introducing additional stochasticity in point selection. An added advantage of this approach is it helps ensure the points selected have geographic validity, i.e., are close enough to actual points because of how we choose the output points based on a dot, which is a very small region. This can been visually confirmed from the generated paths in Figure 7(a), where WildGraph trajectories, much like real trajectories, almost always avoid congregating in the middle of the sea. In contrast, this is not the case with some benchmark methods, including competitive methods such as WildGEN.
(a) Regions of different sizes can be observed
(b) Zooming shows a sample distribution of dots on different regions
(c) Further zooming shows dots as small regions
Figure 5. The breakdown of the spatial domain into regions (blue), and dots (red) projected over a heatmap
@@ 5.1. Baselines
  • GAN
  • Transformer
  • Levy Flight
@@ 5.1.2. WildGEN
  • Minimum Bounding Region: drops any trajectories that venture out of a minimum bounding polygon of the known points.
5.1.3. Generative Adversarial Network (GAN)
GAN is a model consisting of two neural networks, the generator and the discriminator, which engage in a minimax game. The generator aims to produce realistic synthetic data samples from random noise, while the discriminator learns to distinguish between real and generated samples. Through adversarial training, GANs progressively improve both the generator’s ability to create realistic data and the discriminator’s proficiency in discerning real from fake. This adversarial process drives the generator to generate increasingly realistic samples (Goodfellow et al., 2020).
@@ 5.1.4. Transformer
Transformers are networks that have the capacity to auto-regressively generate sequences of tokens by attending to previous tokens. The HNG facilitates the use of H3 word-based language transformers, such that a trajectory is a sentence that consists of $m$ words, each of which is an H3 address.
5.2. Levy Flight
Levy Flight (zaburdaev2015levy) is a stochastic process that models movement using a Cauchy distribution to generate trajectory points. These points are determined by step length, variance, and angular standard deviation, which we measure using the real trajectory samples. The algorithm utilizes these parameters to generate trajectories that attempt to capture and reproduce the movement patterns.
@@ 5.3. Evaluation Metrics
  3. (3)
    (Coverage) The percentage of samples in the held-out set that are closest to any generated sample, for each path similarity metric.
Table 1. Test Results for the geese and stork datasets using WildGraph and other benchmarks using different measurements
Dataset Method Hausdorff DTW FDE SatCLIP r-Coefficient $chi^{2}$
value cov value cov value cov value cov
Geese VAE 7.38 0.56 56.85 0.66 3.67 0.82 0.57 0.53 0.47 1672
GAN 25.53 0.33 194.74 0.25 8.62 0.4 0.44 0.31 0.30 7490
WildGEN 7.15 0.71 48.94 0.82 3.24 0.86 0.6 0.8 0.80 695
WildGraph 9.63 0.93 71.08 0.92 2.93 0.81 0.57 0.89 0.80 628
Transformer 20.82 0.88 216.08 0.78 13.58 0.78 0.40 0.67 0.48 1636
Levy Flight 31.96 0.09 332.14 0.12 29.29 0.08 0.26 0.29 -0.09 9706
Stork VAE 8 0.71 43.3 0.68 4.22 0.65 0.58 0.74 0.56 666
GAN 14.77 0.23 59.95 0.21 3.9 0.42 0.56 0.35 0.62 1491
WildGEN 5.66 0.68 29.26 0.68 3.22 0.68 0.63 0.69 0.77 743
WildGraph 6.61 0.84 35.98 0.81 3.6 0.74 0.59 0.83 0.78 438
Transformer 16.23 0.65 87.19 0.66 8.95 0.58 0.53 0.65 0.56 642
Levy Flight 25.75 0.21 202.21 0.09 31.27 0.11 0.37 0.17 0.09 4492
In order to measure Path Similarity, we use three distance measures: Hausdorff Distance, which measures the steepest gap, Dynamic Time Warping (DTW), which measures the spatial similarity over the whole path, and Final Displacement Error (FDE), which we define as the gap between the final positions of a generated trajectory and the nearest real trajectory: $\text{FDE}=\|p_{t}^{m}-\hat{p}_{s}^{m}\|$.
In addition, we also utilize a satellite imagery-based method, SatCLIP, to measure feature similarity. SatCLIP generates a spatial embedding that summarizes the geographic representation of trajectories, which can then be compared using cosine similarity. The choice of SatCLIP is motivated by its capacity to embed remote geo-locations where other embedding methods, such as hex2vec (Woźniak and Szymański, 2021) were found to provide limited or no data.
On the other hand, we measure Likeness using Pearson Correlation Coefficient (r-Coeff) and Chi-squared ($chi^{2}$) similarity.
For each generated trajectory, we compare it to every trajectory from the test set and use the closest match. The reported distance metric for the experiment is the mean of these values over a fixed number of generated trajectories.
Finally, coverage (cov) measures how representative the generated trajectories are of the test set of real trajectories. This measure indicates what percentage of trajectories in the held-out-set were closest to a generated trajectory. For example, given generated trajectory $S_{i}$, if the best Hausdorff or DTW measure was observed when compared to a real trajectory $T_{j}$, we add $T_{j}$ in the coverage list. This measure quantifies the fraction of trajectories added in the coverage list for each distance metric.
@@ 5.4. Datasets
Table 2. Summary of Geese and Stork Datasets
Dataset Clipping Window Length (days) Sample Size
Western Palearctic greater white-fronted geese (Kölzsch et al., 2019) (Geese) Mar 1 - Sep 1 185 60
White Stork Ciconia ciconia (Carlson et al., 2021) (Stork) Aug 15 - Nov 1 79 77
Table 2 lists the two datasets of migratory birds utilized for testing, Both datasets were obtained from Movebank [^2] https://www.movebank.org and were processed using Moveapps (Kölzsch et al., 2022) to limit observations to one per day for each subject. For both datasets, any subject that had significant missing readings was dropped. Subjects that were tracked over multiple years were split into separate subjects. Simple data-interpolation was utilized to fill in any remaining isolated gaps in the trajectories.
@@ 5.5. Experimental Settings
Preliminary tests were performed to evaluate the convergence properties of each method. We used a fixed learning rate of $1\times 10^{-3}$ with the Adam optimizer. For the GAN experiment, additional decay rate parameters were used to induce convergence.
Each experiment used 5-fold cross-validation with a different held-out set to generalize given the small data nature of the problem. All reported values are the average of five identical experiments. This choice was made to normalize the randomness of the generation process and report more consistent results.
The experiments were conducted utilizing a single 80GB NVIDIA A100 GPU paired with 8 CPUs, operating on Python 3.8, PyTorch 2.1.1, and CUDA 11.8.
@@ 5.5.1. Distributional Settings
To measure the likeness (r-Coeff and $chi^{2}$), we utilize the discrete cluster distributions between the generated points and the test set. To create the clusters, the optimal cluster size for K-means was pre-calculated and used as input in the distributional likeness measurements. The cluster sizes were determined using a Silhouette plot and were selected visually. Specifically, the cluster sizes for the geese and stork datasets are 16 and 14, respectively.
5.5.2. Split Threshold
As discussed in section 4.1, here, we analyze the effect of the split threshold on generalization. Figure 6 depicts how coverage values and r-Coeff are affected by different values of the split threshold.
Figure 6. Coverage as a function of split threshold
For our experiments, we use a split threshold (r) value of 1 which was experimentally validated to provide high coverage, while providing high r-Coeff.
In the ablation study, we explore how the results change as a function of the split threshold.